Given that T = 28.3 kN and W = 25 kN, determine the magnitude and sense of t he moments about point B of the following: a. the force T; b. the force W; c. forces T and W combined.


1
Expert's answer
2021-06-08T09:34:34-0400

Gives

T=28.3kN

W=25N




Moment about point P

(a)


Now

Take components

"F=Tcos\theta"

"F=28.3cos45\u00b0"

F= 20.011 KN

d=20 m

Now torque

"\tau=Tcos45\times20"


"\tau =400.22KN-m"


(b)

Point B corresponding

d=16m

Force of W

W=F=25kN

Torque

"\tau =-16\times25 kN"


"\tau=-400kN- m"


(C)

Net(W and T combination)



"\tau_{net}=400.22-400 =0.22KN-m"